OpenStudy (ttop0816):

x + y = 12 x - y = 10 What is the value of the x-determinant for the system shown? -22 -12 <-- i chose this! -2

OpenStudy (kc_kennylau):

\[Dx=\left|\begin{array}{cc}12&1\\10&-1\end{array}\right|=12\times-1-1\times10=-12-10=-22\]

OpenStudy (ttop0816):

@kc_kennylau oh thank you so much!

OpenStudy (kc_kennylau):

@ttop0816 no problem at all :)

OpenStudy (ttop0816):

@kc_kennylau yay im your 100th fan! congrats (:

OpenStudy (kc_kennylau):

wow :DDDDDD

OpenStudy (anonymous):

x+y=12 x-y=10 x=12-y 12-y-y=10 -2y+12=10 -2y=10-12 -2y=-2 y=1 x=12-y x=12-1 x=11 S={11,1}

OpenStudy (loser66):

@mire but the question is about x determinant , not x. They apply Cramer's rule to have it and they are correct.